Instrument Log 

 

LIRAD C1      GIF Trailer     Lamont, OK 

 

Date

Technician

Comments

 

 

Instrument pulled and returned to mentor.

9/10/01

MR

After the work on the GIF trailer UPS was completed the LIRAD was restarted but failed to reboot properly.  After several attempts to reboot the computer failed the mentor (Richard Austin) was called and a message was left.  On 9/10/01 the mentor was contacted and some T/S was performed over the phone with him.  It was found that the ARM2 file on the disk which contains the operating program for the LIRAD instrument seemed to be corrupted again in the same way it was the last time this instrument was down.  The mentor wants to contact the Australia mentor and confirm with them about shipping the instrument back.  The instrument was shut down until the mentor contacts us regarding shipping instructions and procedure.

9/7/01

MR

Shut the instrument down at 1418 GMT due to work being performed on the GIF trailer UPS.  Will restart as soon as work is complete.

9/6/01

MR

The DOS computer and the FEP computer where locked up.  Rebooted the computers at 1423 GMT.  System came back and started working at 1425 GMT.  RTS at 1438 GMT.

9/5/01

MR

The instrument mentor wants the computers rebooted due to them locking up while he was logged onto them.  T/S found both computers locked up.  Rebooted both computers at 2030 GMT.

9/5/01

MR

Both the DOS and the windows computer were locked up and would not respond to keyboard or mouse control.  Rebooted both computers.  The windows computer restarted ok.  The DOS computer had to be shut down and restarted again due to it not rebooting the first time.  After the second time the computer restarted.  RTS at 1534 GMT.

8/28/01

MR

The instrument mentor wants to check the hard drive to see if it is spinning up. While talking to the mentor on the phone the hard drive was observed to not be working.  Reseated the connections on the hard drive and the hard drive started working but the program would hang up while trying to reboot.  The mentor said he would contact me on further instructions.  I decided to try and run scan disk on the hard drive and found that the FAT had been corrupted.  The FAT can be repaired but the hard drive has to be removed and installed on a machine that has a floppy disk drive.  Removed the hard drive and installed it on a machine with a floppy drive.  Repaired the FAT with Norton Utilities and found that the hard drive also had 11 bad sectors on the disk.  Had the program repair and mark the bad sectors on the disk.  On 8/29/01 I reinstalled the hard drive in the LIRAD and it seemed to reboot and start to run.  Notified the mentor that the instrument was working.  RTS at 1345 GMT on 8/29/01.

6/26/01

MR

The instrument mirror is dirty and needs to be cleaned. Parked the mirror in the up position.  Cleaned the mirror by spraying it with distilled water and then blowing it dry with compressed air.  The mirror was slightly cleaner but still looked discolored.  RTS at 2015 GMT.

6/13/01

JEM

The hatch for the LIRAD was closed, but according to all indicators it should have been in the open position @1515 GMT. A pin was found lying on the ground outside near where the hatch is. That pin had broken off of the hatch arm. Tried to locate another pin similar to the one that had broken, but none were found. Found a hex bolt slightly longer than the pin and placed it through the hatch arm and the hydraulic lift arm @1530 GMT. Placed two nuts on the end of the bolt to ensure that it would not slip out. Opened and closed the hatch several time to ensure that is would not get hung up on the bolt or something else, which it did not. RTS @1545 GMT.

6/11/01

CAM

Richard Austin on site today through Wednesday (6/13) making repairs to the LIRAD computer. A new hard drive was installed but the system crashed again soon thereafter. Further T/S found that some of the RAM was faulty. Replaced the RAM chips and system returned operational. Site Ops received training on how to disassemble the instrument for hard drive replacement.

5/1/01

MR

The DOS machine display is blank and the windows machine is locked up. Rebooted the windows machine and the DOS machine at 1538 GMT. The windows machine came back up OK. The DOS machine came up with "Drive not ready ERROR". Powered cycled the machine but the same error came up again. Sent an E-mail to Richard Austin about the problem. Will await his reply. Shut the DOS machine down at 1542 GMT.

4/30/01

MR

The DOS computer was blank and the FEP computer had no display. The FEP computer had a daily report on the screen with today's report. It looked as if the mentor was working the instrument through the network (the display was changing and different windows where popping up). Will check instrument later today to confirm if it is working properly.

4/24/01

JT

The LIRAD DOS computer experienced a hard drive failure. Richard Austin performed a replacement of the defective hard drive and returned the instrument to operational service.

3/13/01

TMG

LIRAD computer reporting HDD Control Failure. Instrument mentor would like technician to check HDD cabling. Technician check HDD cabling/power connections and rebooted system. Computer still reporting HDD Control failure. Further T/S of system in progress.

3/12/01

MR

Morning rounds found the LIRAD instrument with the following error; "not ready reading drive d:" Tried to press abort but the system locked up. Rebooted the system and got the following error message; "HDD controller not responding". Called Richard in Colorado and notified him. He said he will notify the mentor in Australia. Awaiting further instructions from mentor.

3/8/01

CAM

Longwave Infrared Radiometer (LIRAD) installed in the GIF trailer by Richard Austin and John Bennett. Verbal instructions provided to Site Ops for routine daily checks and mirror cleaning. Richard Austin (Colorado State Univ.) will serve as the primary contact point.